OpenStudy (anonymous):

Find the expression that makes the equation true. 3xy+6y-2x^2+4+_?_=3x^2-5y-7xy-4x and 4y^2+2x^2+9+_?_= 3y^2+x^2+5xy+6

OpenStudy (anonymous):

x=sqrt(100y^2+300y+20N+96)/10+y+2/5 , x=-sqrt(100y^2+300y+20N+96)/10+y+2/5 x=(-sqrt(21y^2-4N-12)+5y)/2, x=(sqrt(21y^2-4N-12)+5y)/2

OpenStudy (anonymous):

Just compare equivalent terms on each side of the equations, the differences are what u need.

OpenStudy (anonymous):

Yeah, I did that, but what is the blank equal to?

OpenStudy (anonymous):

the blank is simply the expression you are trying to find (the bit that makes the equation correct)

OpenStudy (anonymous):

so what are the answers? I still don't understand.

OpenStudy (anonymous):

For example, in the first equation you have -2x^2 on the left. On the right you have +3x^2 To make those pieces equal you need to add 5x^2 on the left. Do the same for all the separate pieces.